OpenStudy (anonymous):

the correct steps to find the value of 16 to the power of 1 over 4?

OpenStudy (anonymous):

Woludn't you do 16*1/4 because 4 is 1/4 of 16? sorry...I'm not sure

OpenStudy (anonymous):

Let's assume 16^(1/4) = x Then 16 = x^4.. Therefore x=?

OpenStudy (anonymous):

that's the original ?

OpenStudy (alekos):

whats another way of writing 16?

OpenStudy (anonymous):

step a

OpenStudy (alekos):

yep, thats it

OpenStudy (anonymous):

can you help me with a few more plz

OpenStudy (anonymous):

this is the question